supposed 3 cons are tossed once, let Y be the random variable representing the number of heads. contruct the probability distribution and the histogram


1
Expert's answer
2022-04-21T08:27:36-0400

Let's denote HH - heads, TT - tails.


There are 23=8 possible outcomes.

S={HHH,HHT,HTH,THH,

HTT, THT, TTH, TTT}

The possible values of the random variable Y are 0, 1, 2, 3.

Possible outcomes

HHH=3

HHT=2

HTH=2

THH=2

HTT=1

THT=1

TTH=1

TTT=0

We will assume that the probability of getting heads and tails is the same:

p=q=1/2

Construct the probability distribution of the random variable

P(0)=P(3)=1/8

P(1)=P(2)=3/8